The water flowing through a 2.0 cm (inside diameter) pipe flows out through three 1.4 cm pipes. (a) If the flow rates in the three smaller pipes are 25, 19, and 12 L/min, what is the flow rate in the 2.0 cm pipe? (b) What is the ratio of the speed of water in the 2.0 cm pipe to that in the pipe carrying 25 L/min?
Answer #1

here,

the water volume flow rate through the pipes , V1 = 25 L/min

V2 = 19 L/min

V3 = 12 L/min

a)

the flow rate in 2 cm pipe , V = V1 + V2 + V3

V = 25 + 12 + 19 L/min

V = 56 L/min

b)

as the diamter of all three pipes is same

their cross sectional area is also same

the ratio of flow rate

V : V1 = 56 : 25

pi * 2^2 * v : pi * 1.4^2 * v1 = 56 : 25

v : v1 = 224 : 12.75